Chapter 8 Homework Help Save & Exit Submit Geved Check my work 7 Cola Inc. and Soda Co. are two of the largest and most successful beverage companies in the world in terms of the products that they sell and their receivables management practices. To evaluate their ability to collect on credit sales, consider the following rounded amounts reported in their annual reports (amounts in millions). 4.32 points Fiscal Year Ended: Net Sales Accounts Receivable Allowance for Doubtful Accounts Accounts Receivable, Net of Allowance 2015 2813 2015 2014 2013 40.75 S.338 $40 6,417 4,664 3,734 ,428 3, 3,763 3,091 50 46 85 4,385 3,713 3,045 6,2784,5793,669 Required: Print 1 Calculate the receivables turnover ratios and days to collect for Cola Inc and Soda Co. for 2015 and 2014. (Use 365 days In a year. Do not round Intermedlate colculations on Accounts Recelvable Turnover Ratlo. Round your final answers to 1 decimal place Use final rounded answers from Accounts Recelvable Turnover Ratlo for Days to Collect ratlo calculatlon.) Cola Inc Soda Co. Cola Inc Soda Co. Receivables Turnover Ratio Days to collect 2 Which of the companies is quicker to convert its receivables into cash? O Soda Co. O Cola Inc.
Answer #1

Answer

  • All working forms part of the answer
  • Workings, Company wise:

Cola Inc.

Working

2015

2014

A

Net Sales

$                32,719

$               28,190

B

Beginning Accounts receivables, net

$                   3,713

$                 3,045

C

Ending Accounts receivables, net

$                   4,385

$                 3,713

D = (B+C) /2

Average Accounts receivables

$                   4,049

$                 3,379

E = A/D

Accounts receivables Turnover Ratio

8.1

8.3

F = 365 days/'E'

Days to collect ratio

45.1

44.0

Soda Co.

Working

2015

2014

A

Net Sales

$                     55,338

$         40,732

B

Beginning Accounts receivables, net

$                        4,579

$            3,669

C

Ending Accounts receivables, net

$                        6,278

$            4,579

D = (B+C) /2

Average Accounts receivables

$                        5,429

$            4,124

E = A/D

Accounts receivables Turnover Ratio

10.2

9.9

F = 365 days/'E'

Days to collect ratio

35.8

36.9

  • Answers

2015

2014

Cola Inc

Soda Co.

Cola Inc

Soda Co.

Receivables Turnover Ratio

8.1

10.2

8.3

9.9

Days to collect

45.1 or 45

35.8 or 36

44.0

36.9 or 37

  • SODA Co. is quicker to converts its receivables into Cash as it has lower Days to Collect ratio compared to Cola Inc.